OpenStudy (anonymous):

Find two numbers whose sum is 61 and whose difference is 33.

OpenStudy (anonymous):

x+y=61 x-y=33

OpenStudy (anonymous):

x+y=61 x-y=33 x=47 and y=14

OpenStudy (anonymous):

thanks LOVES :)